OpenStudy (anonymous):

how do you write this logarithmic expression as a single logarithm : 5 log 3 + log 4

OpenStudy (anonymous):

log 3^5 + log 4 = log (243*4)

OpenStudy (anonymous):

does t make sense? log (a * b) = log a + log b and n log x= log x^n

OpenStudy (anonymous):

yea it does make sense , thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

bt are u sure it is right?

OpenStudy (anonymous):

because isnt it supossed to be like similar to this problem : log 7 + log 2 which equals log 3.5 because u have to divide 7/ 2 .... right ?

OpenStudy (anonymous):

@akash123

OpenStudy (anonymous):

log (a * b) = log a + log b so log 7 + log 2 = log (7 * 2)= log 14

OpenStudy (anonymous):

log 7 - log 2 = log(7/2) = log 3.5 since log a - log b = log(a/b)

OpenStudy (anonymous):

wait i dont get it ... why did u just change ur answer from multiplying to dividing the problem ?

OpenStudy (anonymous):

log ( a *b)= log a+ log b log(a/b) = log a - logb

OpenStudy (anonymous):

in the beginning, it was a typo but i changed it at that time only

OpenStudy (anonymous):

sorry...now it's fine?

OpenStudy (anonymous):

5 log 3 + log 4= log 3^5 + log 4 = log (243*4)

OpenStudy (anonymous):

wait i still dont get it so u multiply that problem instead of dividing ?

OpenStudy (anonymous):

arent u supposed to divide?

OpenStudy (anonymous):

@akash123

OpenStudy (anonymous):

no...log a + log b = log (a*b)

OpenStudy (anonymous):

oh okay thank you , bt why did u put 3^5 instead of putting 5^3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

since n log x= log (x^n) so 5 log 3= log (3^5) = log 243

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

so how do u solve log 8 - 2 log 6 + log 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

log 8 + log 3 - 2 log 6 = log( 8*3) - log (6^2) = log 24 - log 36 = log(24/36)

OpenStudy (anonymous):

okay i have a couple more questions : how would you write these logarithmic expressions as a single logarithm : 1. 4 log m - log n 2. log 5 - k log 2 3.) log [ base of 6] 5 + log [ base of 6 ] x 4. log { base of 7 ] x + log [base of 7] y - log [base of 7] z

OpenStudy (anonymous):

try 1st Q by ur own ...i'll help u

OpenStudy (anonymous):

okay would the answer for the first one be .... log 4^m + log n = log 4/mn

OpenStudy (anonymous):

4 log m = log m^4...right?

OpenStudy (anonymous):

ye i think so

OpenStudy (anonymous):

4 log m - log n = log m^4/n

OpenStudy (anonymous):

oh okay i get it

OpenStudy (anonymous):

so what is the answer for the ne

OpenStudy (anonymous):

next one *

OpenStudy (anonymous):

?